OpenStudy (anonymous):

An equation of the line that passes through the point (0,-1) and whose slope is 2 is A) y=2x+2 B) y=-2x-1 C) y=-x+2 D) y=2x-1

OpenStudy (lieutenantgeneral):

A. y=2x+2

OpenStudy (anonymous):

thx

hartnn (hartnn):

its not A

OpenStudy (anonymous):

@hartnn why?

hartnn (hartnn):

\(y-y_1 = m(x-x_1)\) m = 2 x1 = 0, y1 = -1 plug in the values, what u get ?

OpenStudy (anonymous):

y-(-1)=2(x-0)

hartnn (hartnn):

yess so, y+1 = 2x subtracting 1 from both sides will give you your answer :)

OpenStudy (anonymous):

It's either B or D

OpenStudy (anonymous):

is it D?

hartnn (hartnn):

yes:)